OpenStudy (anonymous):

express using a positive exponent n^-7

OpenStudy (anonymous):

(1/n)^7

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\]

OpenStudy (anonymous):

so yes you could write \[(\frac{1}{n})^7\] or just \[\frac{1}{n^7}\]

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

again correct.